OpenStudy (anonymous):

The length of a rectangle is 3cm longer than its width. Its wiidth is x cm. Its perimeter is 54 cm. Write a equtation that represents the sentence.

OpenStudy (anonymous):

please anwser !!!!

OpenStudy (anonymous):

perimeter of rectangle = 2width + 2length 2x + 2(x + 3) = 54 2x + 2x + 6 = 54 4x + 6 = 54

OpenStudy (anonymous):

4x + 6 = 54 However, all the numbers are divisible by 2, so dividing throughout by 2 we get 2x + 3 = 27